# DATA: Domain-And-Time Alignment for High-Quality Feature Fusion in Collaborative Perception

### v1

@@ 3.2.1 Proximal-region Hierarchical Downsampling (PHD)
where $d_{i}(o_{k})$ is the distance from the ego agent $i$ to the $k$-th object and $d_{th}$ is the distance threshold. To maintain the distribution similarity with the original scene after downsampling, we select $N_{\text{proc}}$ objects from ${\cal{O}}^{\text{prox}}_{i}$ for subsequent processing. If $|{\cal{O}}^{\text{prox}}_{i}|\textgreater N_{\mathrm{max}}$, then $N_{\text{proc}}=N_{\mathrm{max}}$ objects are randomly selected from ${\cal{O}}^{\text{prox}}_{i}$, otherwise, all objects in ${\cal{O}}^{\text{prox}}_{i}$ are selected. Herein, ||$\cdot$|| ⋅ | denotes the cardinality of a set and $N_{\mathrm{max}}$ is the predefined maximum number of objects.
@@
~$\tilde{{\cal{R}}}_{i,k}^{\text{in}}=\text{FPS}({\cal{R}}_{i,k}^{\text{in}},\beta_{\text{in}}),\;\;\tilde{{\cal{R}}}_{i,k}^{\text{out}}=\text{FPS}({\cal{R}}_{i,k}^{\text{out}},\beta_{\text{out}}).$ (3)
@@ 3.2.2 Observability-constrained Discriminator (OD)
First, we define the observations of ego agent $i$ as the ego domain, while the observations of $N-1$ collaborative agents are the collaborative domain. For domain discrimination, one collaborative agent is randomly selected for domain alignment. This enables diverse agent combinations that facilitate domain-invariant feature learning.
@@
where $H_{j}^{\text{comp}}$ and $M_{j}^{\text{comp}}$ represent the complemented feature and observability map, respectively, and the indicator function ${\cal{I}}_{\cal{V}}$ equals to 1 for points in $V$ and 0 for the others. To ensure the domain alignment focus on the regions with shared observability of both agents, an observability weighting map $W\in\mathbb{R}^{1\times H\times W}$ is computed as $W=\text{min}(\text{softmax}([M_{i},M_{j}^{\text{comp}}]))$, where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation and all operations are performed along the first dimension. Finally, the domain alignment objective is
@@ 3.4 Instance-focused Feature Aggregation Module (IFAM)
$(\displaystyle H_{a}^{\text{verif}}=\operatorname{Conv}_{1\times 1}($$(W_{a}^{\text{verif}}\odot H_{a}^{\text{fore}}\oplus(\mathbf{1}-W_{a}^{\text{verif}})\odot H_{a}^{\text{enh}})$ (15)
@@ 4.2 Implementation
The backbone employs PointPillar [15] architecture with a grid size of 0.4m×$\times$0.4m. In CDAM, distance threshold $d_{\text{th}}$ = 50 m and $N_{\text{max}}$ = 2 vehicles are downsampled. The two downsampling ratios are $\beta_{\text{in}}$ = 0.6 and $\beta_{\text{out}}$ = 0.8. For PTAM, the multi-window self-supervised training adopts a window size of $l$ = 16. The training procedure comprises three sequential stages: detection model training, PTAM training, and transmission module training, which are presented in Section 2 of Supplementary Material.
@@ 4.3 Quantitative Results
Robustness to Pose Errors. To investigate the effects of different pose noise components on CP performance, experiments are conducted on V2XSET and V2XSIM. Gaussian noise with varying standard deviations is separately applied to collaborator positions and orientations, with localization noise $\sigma_{\text{loacl}}$ ranging from $0.2~{}m$ to $0.6~{}m$ and heading noise $\sigma_{\text{head}}$ ranging from $0.2^{\circ}$ to $1.0^{\circ}$. As shown in Figure 8, the compared methods suffer from significant performance degradation. Their AP70 declines rapidly with increasing noise magnitude, dropping by over 30% and 9% when $\sigma_{\text{local}}$ and $\sigma_{\text{head}}$ reaches $0.6~{}m$ and $1.0^{\circ}$, respectively, on both V2XSIM and V2XSET. This indicates that spatial feature misalignment severely impacts their CP accuracy. DATA demonstrates robust resilience against pose uncertainties on both datasets: on V2XSIM, it maintains 66.53% AP70 at $\sigma_{\text{local}}=0.6~{}\text{m}$ and 75.23% AP70 at $\sigma_{\text{head}}=1.0^{\circ}$; similarly on V2XSET, it achieves 67.22% AP70 at $\sigma_{\text{local}}=0.6~{}\text{m}$ and 71.17% AP70 at $\sigma_{\text{head}}=1.0^{\circ}$. This robustness stems from two complementary aspects. First, the observability-guided domain alignment learns domain-invariant features from consistent observation, extracting essential geometric patterns from noisy observations. Meanwhile, the instance-focused feature enhancement, through foreground-background separation and semantic refinement, strengthens the structural completeness of object representations, ensuring reliable detection under pose perturbations.
@@ 4.4 Ablation Study
Ablation on Proximal Region Threshold. The PHD module aims to balance point clouds density across different spatial regions while preserving scene structure. The results in Table 4 demonstrate that the $50~{}m$ threshold achieves optimal performance (79.94% AP50), representing a good trade-off between density balancing and distribution preservation. Smaller thresholds ($10~{}m,30~{}m$) provide insufficient coverage of proximal objects, while larger thresholds ($70~{}m,90~{}m$) may introduce excessive downsampling that disrupts the original scene structures.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62271142 and U2341206, in part by the Nanjing U35 Strong Foundation Engineering under Grant 4204002501, in part by the Key Research and Development Program of Jiangsu Province under Grant BE2023021 and BE2023011-2.
@@ 2.2.3 Stage 3: Compression Network
For network optimization, the Adam optimizer [13] is adopted across all three stages. In the first stage, the initial learning rate is set to 0.002 with 10$\times$ decay at epochs 15 and 30, and the models are trained for 40 epochs. The second and third stages both employ a fixed learning rate of 0.001 for 10 epochs. A batch size of 2 is used throughout the entire training process. This sequential training approach allows each component to specialize in its intended functionality while maintaining overall system coherence.
@@ 3.2 Qualitative Comparison Under Synchronous Conditions
Figure 13 presents a demanding collaborative perception scenario with a $280~{}m$ range and noise interference from the V2XSET dataset. The scene contains dense traffic conditions with numerous occlusions and point cloud sparsity at extended distances, creating an ideal testing environment for evaluating collaborative perception capabilities. FFNet and Where2comm achieve good perception results in regions where collaborative agents provide reliable information, but exhibit detection inaccuracies and false positives at greater distances due to point cloud sparsity and noise interference. DATA effectively mitigates noise interference and the challenges of sparse long-distance point clouds through learned domain-invariant features and foreground enhancement, achieving robust detection performance across the entire extended range. Figure 12 similarly demonstrates a noisy scenario from V2XSIM with occlusions and sparse point clouds. DATA achieves effective perception performance, while FFNet and Where2comm exhibit missed detections and false positives in regions with sparse point clouds, further validating DATA’s robust and reliable perception capabilities.
@@ 3.1 Problem Formulation and Overall Architecture
$N$ Number of agents in collaborative perception system Problem Formulation
@@ 3.2.2 Observability-constrained Discriminator (OD)
$sp$ One spatial position CDAM-OD
@@ 4. Experiments and General Symbols
[]delimited-[][$\cdot$][ ⋅ ] Concatenation along specified dimension General Operations
||$\cdot$|| ⋅ | Cardinality of a set General Operations

### v2

@@ 3.2.1 Proximal-region Hierarchical Downsampling (PHD)
where $d_{i}(o_{k})$ is the distance from the ego agent $i$ to the $k$-th object and $d_{th}$ is the distance threshold. To maintain the distribution similarity with the original scene after downsampling, we select $N_{\text{proc}}$ objects from ${\cal{O}}^{\text{prox}}_{i}$ for subsequent processing. If $|{\cal{O}}^{\text{prox}}_{i}|\textgreater N_{\mathrm{max}}$, then $N_{\text{proc}}=N_{\mathrm{max}}$ objects are randomly selected from ${\cal{O}}^{\text{prox}}_{i}$, otherwise, all objects in ${\cal{O}}^{\text{prox}}_{i}$ are selected. Herein, $|\cdot|$ denotes the cardinality of a set and $N_{\mathrm{max}}$ is the predefined maximum number of objects.
@@
$\tilde{{\cal{R}}}_{i,k}^{\text{in}}=\text{FPS}({\cal{R}}_{i,k}^{\text{in}},\beta_{\text{in}}),\;\;\tilde{{\cal{R}}}_{i,k}^{\text{out}}=\text{FPS}({\cal{R}}_{i,k}^{\text{out}},\beta_{\text{out}}).$ (3)
@@ 3.2.2 Observability-constrained Discriminator (OD)
First, we define the observations of ego agent $i$ as the ego domain, while the observations of N1N-1 collaborative agents are the collaborative domain. For domain discrimination, one collaborative agent is randomly selected for domain alignment. This enables diverse agent combinations that facilitate domain-invariant feature learning.
@@
where $H_{j}^{\text{comp}}$ and $M_{j}^{\text{comp}}$ represent the complemented feature and observability map, respectively, and the indicator function ${\cal{I}}_{\cal{V}}$ equals to 1 for points in $V$ and 0 for the others. To ensure the domain alignment focus on the regions with shared observability of both agents, an observability weighting map $W\in\mathbb{R}^{1\times H\times W}$ is computed as $W=\text{min}(\text{softmax}([M_{i},M_{j}^{\text{comp}}]))$, where $\cdot$] denotes concatenation and all operations are performed along the first dimension. Finally, the domain alignment objective is
@@ 3.4 Instance-focused Feature Aggregation Module (IFAM)
$H_{a}^{\text{verif}}=\operatorname{Conv}_{1\times 1}($ $(W_{a}^{\text{verif}}\odot H_{a}^{\text{fore}}\oplus(\mathbf{1}-W_{a}^{\text{verif}})\odot H_{a}^{\text{enh}})$ (15)
@@ 4.2 Implementation
The backbone employs PointPillar [15] architecture with a grid size of $\times0.4m$. In CDAM, distance threshold $d_{\text{th}}$ = 50 m and $N_{\text{max}}$ = 2 vehicles are downsampled. The two downsampling ratios are $\beta_{\text{in}}$ = 0.6 and $\beta_{\text{out}}$ = 0.8. For PTAM, the multi-window self-supervised training adopts a window size of $l$ = 16. The training procedure comprises three sequential stages: detection model training, PTAM training, and transmission module training, which are presented in Section 2 of Supplementary Material.
@@ 4.3 Quantitative Results
Robustness to Pose Errors. To investigate the effects of different pose noise components on CP performance, experiments are conducted on V2XSET and V2XSIM. Gaussian noise with varying standard deviations is separately applied to collaborator positions and orientations, with localization noise $\sigma_{\text{loacl}}$ ranging from 0.2m0.2~m to 0.6m0.6~m and heading noise $\sigma_{\text{head}}$ ranging from $0.2^{\circ}$ to $1.0^{\circ}$. As shown in Figure 8, the compared methods suffer from significant performance degradation. Their AP70 declines rapidly with increasing noise magnitude, dropping by over 30% and 9% when $\sigma_{\text{local}}$ and $\sigma_{\text{head}}$ reaches 0.6m0.6~m and $1.0^{\circ}$, respectively, on both V2XSIM and V2XSET. This indicates that spatial feature misalignment severely impacts their CP accuracy. DATA demonstrates robust resilience against pose uncertainties on both datasets: on V2XSIM, it maintains 66.53% AP70 at $\sigma_{\text{local}}=0.6~\text{m}$ and 75.23% AP70 at $\sigma_{\text{head}}=1.0^{\circ}$; similarly on V2XSET, it achieves 67.22% AP70 at $\sigma_{\text{local}}=0.6~\text{m}$ and 71.17% AP70 at $\sigma_{\text{head}}=1.0^{\circ}$. This robustness stems from two complementary aspects. First, the observability-guided domain alignment learns domain-invariant features from consistent observation, extracting essential geometric patterns from noisy observations. Meanwhile, the instance-focused feature enhancement, through foreground-background separation and semantic refinement, strengthens the structural completeness of object representations, ensuring reliable detection under pose perturbations.
@@ 4.4 Ablation Study
Ablation on Proximal Region Threshold. The PHD module aims to balance point clouds density across different spatial regions while preserving scene structure. The results in Table 4 demonstrate that the 50m50~m threshold achieves optimal performance (79.94% AP50), representing a good trade-off between density balancing and distribution preservation. Smaller thresholds (10m,30m10~m,30~m) provide insufficient coverage of proximal objects, while larger thresholds (70m,90m70~m,90~m) may introduce excessive downsampling that disrupts the original scene structures.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China, in part by the Nanjing U35 Strong Foundation Engineering, in part by the Key Research and Development Program of Jiangsu Province.
@@ 2.2.3 Stage 3: Compression Network
[…]
@@ 3.2 Qualitative Comparison Under Synchronous Conditions
Figure 13 presents a demanding collaborative perception scenario with a 280m280~m range and noise interference from the V2XSET dataset. The scene contains dense traffic conditions with numerous occlusions and point cloud sparsity at extended distances, creating an ideal testing environment for evaluating collaborative perception capabilities. FFNet and Where2comm achieve good perception results in regions where collaborative agents provide reliable information, but exhibit detection inaccuracies and false positives at greater distances due to point cloud sparsity and noise interference. DATA effectively mitigates noise interference and the challenges of sparse long-distance point clouds through learned domain-invariant features and foreground enhancement, achieving robust detection performance across the entire extended range. Figure 12 similarly demonstrates a noisy scenario from V2XSIM with occlusions and sparse point clouds. DATA achieves effective perception performance, while FFNet and Where2comm exhibit missed detections and false positives in regions with sparse point clouds, further validating DATA’s robust and reliable perception capabilities.
@@ 3.1 Problem Formulation and Overall Architecture
NN Number of agents in collaborative perception system Problem Formulation
@@ 3.2.2 Observability-constrained Discriminator (OD)
spsp One spatial position CDAM-OD
@@ 4. Experiments and General Symbols
$\cdot$] Concatenation along specified dimension General Operations
$|\cdot|$ Cardinality of a set General Operations